Estimate no. Of collisions per second suffered by a molecule in a sample of hydrogen at STP. The mean free path = 1.3*10^-5

Grade:11

1 Answers

Sumantrak Mukherjee
6 years ago
((8×R×T)÷(π×M))=average velocity(V)At STP→ T=273 K molecular mass of H2=2g/molR=8.3V=√((8×8.3×273)÷(π÷2×1,000)) [mass to be taken in kg]1698.53Mean free path=1.38×10^(-5)cmTime elapsed =mean free path÷average velocityT=1.38×10^(-7)÷1698.530.0008124673×(10^-7)Number of collisions in one second =1÷T1÷(8.124673×10^-11)=12,308,187,664.906≈1.23×10^10